OpenStudy (anonymous):

revolve the region bounded by y = x^2 , y = 1, about the y axis .

OpenStudy (anonymous):

before you say anything, its not possible to do. x = sqrt y, and x = -sqrt y , but i have

OpenStudy (dumbcow):

\[V = \pi \int\limits_{0}^{1}(\sqrt{y})^{2} dy\]

OpenStudy (dumbcow):

what do you mean?

OpenStudy (anonymous):

but what about -1 to 0

OpenStudy (dumbcow):

function is not defined for y<0

OpenStudy (anonymous):

but the region is defined

OpenStudy (anonymous):

ok here is another one like it, revolve y = x^2 and y = x+2 bounded, about y axis

OpenStudy (anonymous):

it exists in the math universe, not in the real world ( you can only revolve one revolution at a time about an axis )

OpenStudy (anonymous):

dumbcow, do you see y = x^2 , y = x+2

OpenStudy (dumbcow):

oh yes i see, sorry i was also assuming it was bounded by x=0

OpenStudy (anonymous):

so does it make sense , you get extra volume, correct ?

OpenStudy (dumbcow):

wait no it would still be same volume if rotating around y-axis the whole region is from y=0 and y=1

OpenStudy (anonymous):

but you have extra volume in the second quadrant

OpenStudy (dumbcow):

hmm ok its double the volume then

OpenStudy (anonymous):

not quite double though

OpenStudy (anonymous):

its a real feather of a problem, lets use shell method , see what we get

OpenStudy (anonymous):

easier than using disc method here

OpenStudy (anonymous):

integral 2pi * x ( x+2 - x^2)

OpenStudy (dumbcow):

ok \[=2 \pi \int\limits_{-1}^{1}x(x ^{2})dx\]

OpenStudy (dumbcow):

oh we are doing the other one

OpenStudy (anonymous):

as i suspected, we get negative volume from -1 to 0

OpenStudy (anonymous):

i think in the shell method, its important that the radius remain positive

OpenStudy (anonymous):

here is a similiar question, revolve region y = (x+2)^2 and y = 1 about y axis, using shell method. you get a negative answer

OpenStudy (dumbcow):

yes but volume can't be neg in the practical sense

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so there are limitations to the revolving region problems

OpenStudy (dumbcow):

you just switch the bounds from neg to positive, because when revolving around axis there is symmetry

OpenStudy (dumbcow):

not in every case though

OpenStudy (anonymous):

what about y = x^2 - 4 and bounded by y = 0, revolve it about x axis

OpenStudy (anonymous):

i think that works, we only get problems with the shell method because the radius is assumed to be positive

OpenStudy (anonymous):

and in the disc method we have radius squared, so its more flexible, but why cant we integrate y = x^2 and y=1 about y axis, using horizontal discs ?